### v1

# Scalable Two-Stage Stochastic Optimal Power Flow via Separable Approximation

Abstract
This paper proposes a Separable Projective Approximation Routine – Optimal Power Flow (SPAR-OPF) framework for solving two-stage stochastic optimization problems in power systems. The framework utilizes a separable piecewise linear approximation of the value function and learns the function based on sample sub-gradient information. We present two formulations to model the learned value function, and compare their effectiveness. Additionally, an efficient statistical method is introduced to assess the quality of the obtained solutions. The effectiveness of the proposed framework is validated using distributed generation siting and sizing problem in three-phase unbalanced power distribution systems as an example. Results show that the framework approximates the value function with over 98% accuracy and provides high-quality solutions with an optimality gap of less than 1%. The framework scales efficiently with system size, generating high-quality solutions in a short time when applied to a 9500-node distribution system with 1200 scenarios, while the extensive formulations and progressive hedging failed to solve the problem.
Index Terms:
Decision-making under uncertainty, power systems planning, stochastic optimal power flow, value function approximation.
I Introduction
The reliable, secure, and economical operation of power systems relies on solving a variety of optimization problems. Relevant applications include economic dispatch (ED), unit commitment (UC), market clearing, reserve planning, system restoration, expansion planning, and the optimal placement of distributed generation [1]. These optimization challenges encompass both transmission and distribution systems, with their complexity depending on factors such as problem type, decision time horizon, and the nature of the decision variables. Traditionally, grid optimization focused on a limited set of credible scenarios, where nonlinearity of power flow and integer variables mainly posed significant challenges. However, the integration of variable energy resources, flexible/controllable loads, and pronounced weather-related uncertainties have increased the complexity of grid operations significantly. Modern grids must now be optimized under a much broader range of scenarios, making optimization under uncertainty a far greater challenge that demands scalable and robust methods.
Optimization under uncertainty has been widely studied in the power systems domain, particularly in relation to various Optimal Power Flow (OPF)-based problems. In this context, the power systems community has explored a range of formulations, including stochastic, robust, distributionally robust, and chance-constrained optimization, as well as their hybrid forms [2]. Many real-world power system challenges, such as UC, ED, proactive dispatch, and investment planning, naturally lend themselves to a two-stage decision-making framework. In this setup, the first stage involves making non-anticipative decisions, while the second stage optimizes operational outcomes once uncertainty is realized. However, for large-scale systems, solving these two-stage problems becomes computationally intractable, particularly in resource-constrained environments [3]. This presents a critical gap in current methods, as the size and complexity of real-world power systems continue to grow. In this paper, we present a novel, computationally efficient, and scalable framework for solving two-stage stochastic optimization problems in power systems. The proposed method leverages a separable piecewise linear approximation of value function to achieve near-optimal solutions, significantly enhancing tractability for large-scale systems while maintaining high solution quality.
I-A Background and Related Work
Traditional approaches for solving two-stage optimization problems include chance-constrained optimization, robust optimization, and stochastic optimization. Chance-constrained optimization requires knowledge of uncertain parameter distributions and generally results in an optimization problem of higher complexity (e.g. non-convex problems, even when the original problem is convex) [3]. Robust optimization tends to be overly conservative, resulting in inefficient resource utilization. In contrast, stochastic optimization directly uses scenario sets as input, considering recourse actions in the second stage to ensure feasibility, thus eliminating the need to model or approximate uncertainty distributions. This motivates us to consider an approach based on stochastic optimization.
Ideally, two-stage stochastic optimization involves solving a deterministic extensive formulation (EF) where multiple copies of the second stage problem are added to the model for each scenario [4]. Solving such an extensive form is computationally intractable, as the size of the decision space increases exponentially with the number of scenarios. This is because the number of possible realizations increases exponentially with the number of uncertain parameters and the system size [5]. However, with the increasing level of uncertainty, it is not advisable to reduce excessively the number of scenarios just for computational purposes as a reasonable number of scenarios need to be considered in order to capture the uncertainty—otherwise, the obtained solutions may be sub-optimal or infeasible [6].
Traditional approaches to tackle scalability challenges include (1) solving the EF multiple times with reduced scenario size, e.g., using a Sample Average Approximation (SAA), (2) using a stage-wise decomposition, e.g., Benders decomposition (BD), and (3) using a scenario-wise decomposition, e.g., progressive hedging (PH) [3]. SAA involves solving the EF problem multiple times while taking a smaller sample size from the set of scenarios to approximate the value function. For example, SAA was used to minimize load consumption in distribution systems through conservation voltage reduction on the IEEE 37 bus system and 2000 scenarios [7]. However, for a realistic system size, the problem poses significant computational challenges even with a relatively small number of scenarios. BD and its variants have been utilized in solving various OPF problems. For instance, BD was used to solve the unit commitment problem for the IEEE 118 bus system considering 10 scenarios [8]. Stochastic decomposition, a variant of BD, has been applied in [9] for solving the economic dispatch problem considering a maximum system size of 96 buses and a maximum number of 100 scenarios. As the number of scenarios grows, the demand for generating valid cuts increases rapidly, making the problem intractable and unsuitable for larger systems with numerous scenarios. It was also observed that BD often leads to long computational times, excessive use of memory resources, poor feasibility and optimality cuts, and slow convergence (tail-off effect) [10]. On the other hand, PH, an application of the augmented Lagrangian relaxation, results in scenario subproblems which are then solved iteratively until convergence criteria are met [11]. PH was used in [12] for proactive dispatch and investment planning, considering the 2000 bus system and 512 scenarios. However, in real-world applications, the default settings for PH often pose challenges [13]. Moreover, due to its decompositional nature, PH is most effective in parallel architectures with multi-core devices, thereby restricting its broader applicability.
Separate from these methods, the optimization community has explored the use of learning-based approaches to solve two-stage stochastic optimization problems. One popular approach includes a separable approximation of a value function where the coordinate functions are approximated using a learning-based technique. A static piecewise linear approximation was proposed in [14] , but this method, being not adaptive, limits the quality of the solution. Later, a Separable Projective Approximation Routine (SPAR) was introduced, which adaptively constructs a sequence of piecewise linear functions using sample sub-gradient information [15]. It proved to be optimal when the value function is separable while also providing a near-optimal solution for the non-separable case, which proved faster than BD [15]. Unlike popular black-box neural network models, these methods maintain the explainability of the learned functions and hence are preferable for power systems, where interpretability of decision-making processes are crucial for safety, reliability, and regulatory compliance.
I-B Contributions
We introduce Separable Projective Approximation Routine - Optimal Power Flow (SPAR-OPF), an efficient framework designed to solve two-stage stochastic optimization problems involving OPF in power systems. Leveraging a SPAR-based approach, our framework scales effectively with system size and can be implemented in resource-constrained environments. To the best of our knowledge, this is the first framework to solve OPF-based two-stage stochastic optimization problems using separable projective approximations of value function. From the application perspective, we consider the following two-stage setting: Stage-1 makes ex-ante decisions based on forecasts, and Stage-2 makes ex-post decisions after the realization of the event/uncertainty. Stage-1 can be thought of as a long-term or operational planning problem, and Stage-2 represents real-time operations. The general assumption is that planning decisions should be optimized considering optimal real-time operations. The computational performance of SPAR-OPF indicates its applicability for solving OPF under uncertainty for realistic power systems. The following are the main contributions of this paper.
  • SPAR-OPF framework for two-stage stochastic optimization: We propose a framework to solve two-stage stochastic OPF problems in power systems based on separable projective approximation of value function. We present two formulations: the lambda method and the epigraph method to model the learned value function and compare their performance.
  • Posterior statistical bounds for optimal solutions: We propose a computationally efficient statistical method that uses a posterior analysis to obtain lower and upper bounds for the optimal objective function value as measures of quality of the obtained solutions.
  • Validation and scalability assessment: We validate and assess the scalability of SPAR-OPF on a two-stage OPF problem optimizing the siting and sizing decisions of PV-based distributed generation (DG) for enhanced and efficient grid operation. The proposed SPAR-OPF approach is compared with approaches using Extensive Formulations (EF) and the Progressive Hedging (PH) algorithm. Finally, the scalability of SPAR-OPF is demonstrated on a 9500-node distribution test system.
II Separable Projective Approximation Routine – Optimal Power Flow (SPAR-OPF) Framework
II-A Problem Formulation
We present SPAR-OPF as a general two-stage stochastic optimization model with optimal planning in the first stage and optimal operation (using OPF) in the second stage. Subsequently, specific details for the framework using DG siting and sizing as an example of a two-stage stochastic OPF problem are presented. The first stage problem is denoted as $\textbf{P}_{\textbf{g1}}$ and is defined in (1). It is important to mention that we do not impose any restrictions on the nature or structure of $\textbf{P}_{\textbf{g1}}$. The objective function consists of a first-stage cost $H(x,y)$ that is dependent only on the first-stage decision variables $x$ and $y$, and the expected recourse cost $Q(x,\xi)$. We obtain $Q(x,\xi)$ by solving the second stage OPF problem denoted by $\textbf{P}_{\textbf{g2}}$ with $z^{\xi}$ as decision variables, as given in (3). OPF has been shown to be exact under various convex relaxations and approximations [16]. We assume such convex problems in the second stage. Here, $x$ is the set of linking variables involved in the second stage for each scenario $\xi$. Note that the constraints in the second stage problem are separated into two types: one dependent on the first stage decision variable $x$, denoted as linking constraints (4), and the rest of the constraints (5)–(6) that do not directly involve the linking variables. Most general two-stage power grid optimization problems can be reformulated to fit the problem structure defined in (1)–(6).
[…]
$(\textbf{P}_{\textbf{g1}}):\quad\min_{x,y}~{}H(x,y)+\mathbb{E}_{\xi}[Q(x,\xi)]$ (1)
[…]
 s.t.    {planning constraints} (2)
$(\textbf{P}_{\textbf{g2}}):\quad Q(x,\xi)=~{}\min_{z^{\xi}}~{}f_{0}(z^{\xi})\hskip 28.90755pt$ (3)
[…]
$f_{i}(z^{\xi})\leq 0\quad\forall\ i\in 1,2,...m$ (5)
[…]
$B^{\xi}z^{\xi}=e^{\xi}\quad$ (6)
Proposition 1.
If the second stage problem is convex with continuous recourse variables, $Q(x,\xi)$ is a piecewise linear convex function in the linking variable $x$.
Proof.
[…]
$\begin{gathered}L(x,\xi,z^{\xi},\alpha,\beta,\nu)=f_{0}(z^{\xi})+\alpha^{T}(A^{\xi}z^{\xi}-x)+\\ \sum_{i=1}^{m}\beta_{i}f_{i}(z^{\xi})+\nu^{T}(B^{\xi}z^{\xi}-e^{\xi})\end{gathered}$ (7)
The Lagrangian dual function $q(x,\xi,\alpha,\beta,\nu)$ and the associated dual problem optimal value $q^{D}(x,\xi)$ are then obtained as
[…]
$\begin{gathered}q(x,\xi,\alpha,\beta,\nu)=\inf_{z^{\xi}}~{}f_{0}(z^{\xi})+\alpha^{T}(A^{\xi}z^{\xi}-x)+\quad\quad\\ \sum_{i=1}^{m}\beta_{i}f_{i}(z^{\xi})+\nu^{T}(B^{\xi}z^{\xi}-e^{\xi})\quad\text{\rm and}\end{gathered}$ (8)
[…]
$\begin{gathered}q^{D}(x,\xi)=\max_{\alpha,\beta,\nu}~{}q(x,\xi,\alpha,\beta,\nu)\end{gathered}$ (9)
The Lagrangian dual problem (9) is convex. With the assumption that Slater’s condition is satisfied for the inequalities in (4) and (5), we get by strong duality that $Q(x,\xi)=q^{D}(x,\xi)$. From the expression for $q(x,\xi,\alpha,\beta,\nu)$ in (8), it follows that $q^{D}(x,\xi)$ is the maximum of affine functions of the linking variable $x$ through the term $-\alpha^{T}x$. Hence, $Q(x,\xi)$ is a piecewise linear convex function of $x$. ∎
Remark 1.
Under the assumption that the probability distribution of uncertain parameters is independent of decisions made and that Proposition 1 holds, $\mathbb{E_{\xi}}[Q(x,\xi)]$ is a piecewise linear convex function as it is the linear combination (coefficients are probabilities of scenarios) of piecewise linear convex functions.
II-B Value Function Approximation (VFA):
The value function $\mathbb{E_{\xi}}[Q(x,\xi)]$ in the objective function (1) of the first stage problem $\textbf{P}_{\textbf{g1}}$ is a function of linking variables $x$ and scenarios $\xi$ realized in the second stage. We write the value function as a summation of coordinate functions $g_{i}(x_{i})$ [^2] Coordinate function $g_{i}(x_{i})$ is a function of coordinate variable $x_{i}$ as given in (10) where $\mathcal{I}$ is the set of coordinate functions.
[…]
$G(x)=\mathbb{E}_{\xi}[Q(x,\xi)]=\sum_{i=1}^{|\mathcal{I}|}g_{i}(x_{i})$ (10)
Let us assume $g_{i}(x_{i})$ is a piecewise linear convex function with integer breakpoints in $x_{i}$ dimension as represented by the blue curve in Fig. 1. Let $m_{i}^{l}$ be its slope between points $l\ \in\mathcal{L}_{i}$ and $l+1$ as given in (11). The value of the function at any point $x_{i}$ can be obtained using (12) where $u$ is an integer point.
$m_{i}^{l}=g_{i}(l+1)\ -\ g_{i}(l)$ (11)
[…]
$\begin{gathered}g_{i}(x_{i})=\begin{cases}g_{i}(0)-\sum_{l=u}^{-1}m_{i}^{l}+m_{i}^{u}(x_{i}-u)\,\text{ for }x_{i}<0\\ g_{i}(0)+\sum_{l=0}^{u-1}m_{i}^{l}+m_{i}^{u}(x_{i}-u)\,\text{ for }x_{i}\geq 0\end{cases}\\ \text{where}\ u\leq x_{i}<u+1.\end{gathered}$ (12)
Figure 1: Piecewise linear convex functions with integer break points.
It is to be noted that, evaluating this function requires its value at origin, $g_{i}(0)$, which may not be available a priori. The decision points, however, are not affected by this constant term, and hence, $g_{i}(0)=0$ can be assumed. This shifts the function towards the origin as represented by the red curve in Fig. 1.
Once the coordinate functions are learned, a formulation for (10) is constructed, and the first stage problem $\textbf{P}_{\textbf{g1}}$ can then be solved by considering only the first stage feasible set. In this regard, we propose two formulations. First, we introduce binary variables to approximate the function using a well-known polyhedral approach to model piecewise linear functions [17], which we refer to as the lambda method $(\textbf{L}_{\textbf{1}})$. However, introducing a large number of binary variables can significantly increase computational complexity. To mitigate this challenge, we propose a second approach that we refer to as the epigraph method $(\textbf{E}_{\textbf{1}})$, which leverages the convexity of the value function by incorporating a continuous epigraph variable for each coordinate function.
II-B1 Lambda Method
Once the slope of the coordinate function $g_{i}$ is learned, the function can be computed using (12). We introduce binary variables $\lambda_{i}^{l}$ associated with each decision point $l\in\mathcal{L}$ of the coordinate function $g_{i}$ to approximate (10) as specified in the objective function in (13). Here, $g_{i}^{l}$ represents the value of the function $g_{i}$ at point $l$. Eq. (14) ensures that each coordinate function has a single decision point. The corresponding linking decision variable $x_{i}$ is then given by (15).
[…]
$(\textbf{L}_{\textbf{1}}):\quad\min~{}H(x,y)+\sum_{l\in\mathcal{L}_{i}}\sum_{i\in\mathcal{I}}g_{i}^{l}\lambda_{i}^{l}$ (13)
[…]
$\sum_{l\in\mathcal{L}}\lambda_{i}^{l}l=x_{i}~{}~{}\forall\ i\in\mathcal{I}$ (15)
[…]
planning constraints (2)
[…]
$\lambda_{i}^{l}\in\{0,1\},~{}\forall\ l\in\mathcal{L}_{i},\forall\ i\in\mathcal{I}$
Remark 2.
Constraint (15) ensures that $x_{i}$ is integral in any feasible solution even when it is modeled as a continuous variable, since $l\in\mathcal{L}_{i}$ is an integer parameter and $\lambda_{i}^{l}\in\{0,1\}$.
II-B2 Epigraph method
Since the learned function is piecewise linear and convex with integer breakpoints, minimizing the function (1) after replacing the expectation term with (10) is equivalent to minimizing the function (16) with additional constraints (17). The function $g_{i}$ is approximated by the affine function using slope information $m_{i}^{l}$ at each integer breakpoint $l$. This function is then upper bounded by the epigraph variable $t_{i}$.
[…]
$(\textbf{E}_{\textbf{1}}):\quad\min\ H(x,y)+\sum_{i\in\mathcal{I}}\,t_{i}$ (16)
[…]
planning constraints (2)
II-C Separable Projective Approximation Routine (SPAR)
In this section, we present the SPAR algorithm from [15] for learning the coordinate functions $g_{i}$. Any term with the superscript $(k)$ refers to its value at iteration $k$. For simplicity, we represent $g_{i}(x_{i})$ as $g(x)$ and $m_{i}^{l}$ as $m_{l}$, which is the slope of function $g(x)$ between points $l$ and $l+1$. The overall goal is to learn and update the slopes iteratively such that the sequence $\{m^{(k)}\}$ converges to the true slope $\overline{m}$. In other words, the function $\{g^{(k)}\}$ constructed from the learned slopes using (12) is the functional approximation of the true function $g$. Algorithm 1 specifies how to learn the slopes of the coordinate function $g$. First, the slope vector $m^{(1)}$ is initialized to the zero vector and the maximum number of iterations is chosen as $k^{\max}$. Then, the first stage decision point $l^{(k)}$ is obtained by solving the first stage problem $\textbf{P}_{\textbf{g1}}$ after replacing the expected function by approximated function $G^{(k)}$ using (10). Notice that in initial iterations, exploration can be employed by randomly choosing decision points, rather than exploiting the optimal decisions in every iteration. After this, a random variable $\gamma^{(k)}$ is observed, which is $H_{k}$ measurable [^2] It means this value is based on the information available up to $k$ iterations, and its expectation is equal to the true slope $\overline{m}_{l^{(k)}}$. As discussed in [15], $\gamma^{(k)}$ satisfies this condition when Lagrange multipliers are taken after solving the second-stage problem for the given first-stage decision $l^{(k)}$. Next, the slope is updated at point $l^{(k)}$ as shown in step 5. Step size $\{\alpha^{(k)}\}$ is also $H_{k}$ measurable and assumed to satisfy $\sum_{k=1}^{\infty}\alpha^{(k)}=\infty$ and $\sum_{k=1}^{\infty}(\alpha^{(k)})^{2}<\infty$. The updated slope vector $n^{(k)}$ is then projected on a convex set $\mathcal{M}$ given by (19) to give slope vector $m^{(k+1)}$ for the next iteration to ensure the convexity of the learned function. Finally, the function $G^{(k)}$ is updated to give $G^{(k+1)}$ and the process continues until $k^{\max}$ iterations are reached or the termination criteria are met. One of the termination criteria considered in this paper is that the algorithm terminates when the moving average of the objective value at the current iteration falls within an $\epsilon$-radius of the moving average of the previous iteration.
Algorithm 1 Separable Projective Approximation Routine (SPAR)
1:  Initialize: $k=1,m^{(1)}=\textbf{0}^{L}$
2:  for $k=1,2,\dots,k^{\max}$ do
3:     Get $l^{(k)}\in[\underline{L},\overline{L}]$ after solving $\textbf{P}_{\textbf{g1}}$
4:     Observe scenario and get random variable $\gamma^{(k)}$ after solving $\textbf{P}_{\textbf{g2}}$ such that,$\mathbb{E}(\gamma^{(k)}|l^{(1)},\dots,l^{(k)},\gamma^{(1)},...,\gamma^{(k-1)})=\overline{m}_{l^{(k)}}$
5:     Update slope as:
$n_{x}^{(k)}=\begin{cases}(1-\alpha^{(k)})m_{x}^{(k)}+\alpha^{(k)}\gamma^{(k)}\ \textbf{if}\ x=l^{(k)},\\ m_{x}^{(k)}\ \textbf{if}\ x\neq l^{(k)}\end{cases}$
6:     Project $n^{(k)}\ \text{to convex set}\ \mathcal{M}\ \text{to yield}\ m^{(k+1)}$
7:     Check the termination criteria
8:  end for
II-D Projection Algorithm
[…]
$\min_{m}\ \frac{1}{2}\|m-n\|$ (18)
[…]
Taking $\beta_{l}$ as dual multipliers of (19), the optimality conditions for this problem can be written as,
[…]
$m_{l}=n_{l}+\beta_{l-1}-\beta_{l}\quad\forall\ l=\underline{L}+1,\underline{L}+2,\dots,\overline{L}$ (20)
[…]
$\beta_{l}(m_{l}-m_{l+1})=0\quad\forall\ l=\underline{L},\underline{L}+1,\dots,\overline{L}$ (21)
[…]
$m_{p-1}<m_{p}=...=d=...=m_{q}<m_{q+1}$ (22)
Adding (20) from $l=p$ to $l=q$ and using (22), we obtain
[…]
$\sum_{l=p}^{q}m_{l}=\sum_{l=p}^{q}n_{l}+\beta_{p-1}-\beta_{q}$ (23)
[…]
$(q-p+1)d=\sum_{l=p}^{q}n_{l}+\beta_{p-1}-\beta_{q}$ (24)
From (21) and (22), $\beta_{p-1}=0,\beta_{q}=0$, (24) becomes,
[…]
$d=\frac{1}{(q-p+1)}\sum_{l=p}^{q}n_{l}$ (25)
In Algorithm 1, the slope vector is updated at a single coordinate, say $l^{(k)}$ at iteration $k$. Considering $n_{l}^{(k)}$ as a slope at point $l$ at iteration $k$, the updated slope at the current chosen point $l^{(k)}$ would be $n_{l^{(k)}}^{(k)}$. Since the slope is updated at only one point, the following three conditions can occur.
  • Condition 1: $n_{l^{(k)}-1}^{(k)}\leq n_{l^{(k)}}^{(k)}\leq n_{l^{(k)}+1}^{(k)}$
    The updated slope maintains the convexity. Hence, no further update is required.
  • Condition 2: $n_{l^{(k)}-1}^{(k)}>n_{l^{(k)}}^{(k)}$
[…]
    $n_{\eta-1}^{(k)}\leq\frac{1}{l^{(k)}-\eta+1}\sum_{l=\eta}^{l^{(k)}}n_{l}^{(k)}$
    and assign, $m_{l}^{(k+1)}=d\quad\forall\ l=\eta,~{}\eta+1,\dots,l^{(k)}$, where $d$ is obtained from (25) with $p=\eta\ \text{and}\ q=l^{(k)}$. If no $\eta$ is found, assign $\eta=\underline{L}$.
  • Condition 3: $n_{l^{(k)}}^{(k)}>n_{l^{(k)}+1}^{(k)}$
[…]
    $n_{\eta+1}^{(k)}\geq\frac{1}{\eta-l^{(k)}+1}\sum_{l=l^{(k)}}^{\eta}n_{l}^{(k)}$
    and assign, $m_{l}^{(k+1)}=d\quad\forall\ l=l^{k},l^{k}+1,\dots,\eta$ where $d$ is obtained from (25) with $p=l^{k}\ \text{and}\ q=\eta$. If no $\eta$ is found, assign $\eta=\overline{L}$. The Lagrange multipliers corresponding to the above solutions exactly matches with that of (20) and (21[15].
II-E Statistical method for lower and upper bound
Since the proposed framework uses sample subgradient information to approximate the function, we provide a statistical method for evaluating the quality of the obtained solution through a posteriori analysis, utilizing the information available during the process of solving the stochastic optimization problem. We denote the objective function $H(x,y)+\mathbb{E}_{\xi}[Q(x,\xi)]$ in (1) as $F(x,y)$ and let $F^{*}$ be its optimal value. First, we provide the statistical upper bound, i.e., an estimate of the upper bound and its confidence interval at a chosen level. Let $(\hat{x},\hat{y})$ be the decisions obtained after solving (1) where the function is approximated using SPAR-OPF taking batches of $N^{\prime}$ samples from scenarios set $\Xi$. Consider that $p_{\xi}$ is the probability of scenario $\xi$ and $h_{|\Xi|}$ is the value of (1) where expected function is calculated by simulating all scenarios independently in the second stage $\textbf{P}_{\textbf{g2}}$ fixing the first stage decision $(\hat{x},\hat{y})$ as given in (26). This gives the upper bound for $F^{*}$. Taking expectation in the left side of (26) also gives the upper bound of $F^{*}$ as given in (27) where (1) is solved $M$ times taking batches of scenarios. The unbiased estimator, $\bar{h}_{|\Xi|}$, of $\mathbb{E}$($h_{|\Xi|}$) and variance $\sigma_{M}^{2}$ are given in (28) and (29), respectively. The confidence interval for the upper bound $UB_{M}$ with an approximated (1-$\alpha$) confidence is given in (30), which is justified by the central limit theorem with critical value $z_{\alpha}=f^{-1}(1-\alpha)$ where $f(z)$ is the cumulative distribution function of the standard normal distribution. Note that if the value of $M$ is low, the $t$-distribution should be used.
[…]
$h_{|\Xi|}=c^{T}\hat{x}+d^{T}\hat{y}+\sum_{i\in 1}^{|\Xi|}p_{\xi}Q(\hat{x},\xi_{i})\geq F^{*}$ (26)
[…]
$\mathbb{E}(h_{|\Xi|})\geq F^{*}$ (27)
[…]
$\bar{h}_{|\Xi|}=\sum_{i\in 1}^{M}h_{|\Xi|}^{i}$ (28)
[…]
$\sigma_{M}^{2}=\frac{1}{M(M-1)}\sum_{i=1}^{M}[\bar{h}_{|\Xi|}-h_{|\Xi|}^{i}]^{2}$ (29)
[…]
$UB_{M}=\bar{h}_{|\Xi|}\pm z_{\alpha}\sigma_{M}$ (30)
The lower bound for $F^{*}$ can be obtained by leveraging the sample average approximation. The sample average approximation of $F_{N^{{}^{\prime}}}(x,y)$ with $N^{{}^{\prime}}$ samples is an unbiased estimator of $F(x,y)$ as given in left equality in (31). If $p_{N^{{}^{\prime}}}$ is an optimal value of (1) for $N^{{}^{\prime}}$ samples, then $\mathbb{E}(p_{N^{{}^{\prime}}})$ is a lower bound for $F(x,y)$ as given in (31). Taking minimum of $F(x,y)$ and using (31), we obtain (32). The unbiased estimator $\bar{p}_{N^{{}^{\prime}}}$ of $\mathbb{E}(p_{N^{{}^{\prime}}})$ and the variance are given in ( 33) and (34), respectively, where problem (1) is solved $M$ times taking $N^{{}^{\prime}}$ samples. The confidence interval for the Lower bound $LB_{N^{{}^{\prime}}}$ of $F^{*}$ with an approximated (1-$\alpha$) confidence is given by (35).
[…]
$F(x,y)=\mathbb{E}[{{F}_{N^{\prime}}(x,y)}]\geq\mathbb{E}[{\min_{x,y}{F}_{N^{\prime}}(x,y)]}=\mathbb{E}({p_{N^{\prime}}})$ (31)
[…]
$\min_{x,y}F(x,y)=F^{*}\geq\mathbb{E}{(p_{N^{\prime}})}$ (32)
[…]
$\bar{p}_{N^{{}^{\prime}}}=\sum_{i\in 1}^{M}p_{N^{{}^{\prime}}}^{i}$ (33)
[…]
$\sigma_{N^{{}^{\prime}}}^{2}=\frac{1}{M(M-1)}\sum_{i=1}^{M}[\bar{p}_{N^{{}^{\prime}}}-p_{N^{{}^{\prime}}}^{i}]$ (34)
[…]
$LB_{N^{{}^{\prime}}}=\bar{p}_{N^{{}^{\prime}}}\pm z_{\alpha}\sigma_{N^{{}^{\prime}}}$ (35)
Remark 3.
Since SPAR-OPF solves the problem by approximating the value function, the approximation quality can be measured as the iterations proceed. After approximating the function, either $\textbf{L}_{\textbf{1}}$ or $\textbf{E}_{\textbf{1}}$ is solved to get first stage decisions $(x^{(k)},y^{(k)})$, and then $H(x^{(k)},y^{(k)})+\mathbb{E}_{\xi}[Q(x^{(k)},\xi)]$ is computed to measure quality of the approximated function.
III Stochastic DG Siting and Sizing Problem
The proposed framework is implemented to solve the optimal DG siting and sizing problem in distribution systems. The distribution system is modeled as a graph $\mathcal{G(B,E)}$ where the node $i\in\mathcal{B}$ represent the bus and the edge $ij\in\mathcal{E}$ represents the line that connects nodes $i$ and $j$. We consider a three-phase unbalanced linear power flow model to represent the power flow physics for the distribution system [18]. The decisions variables and constraints associated with the two-stage problem are described below.
III-A First Stage
The objective function in the first stage problem D1 is given by (36), which minimizes the expected recourse cost $Q(x,\xi)$. The first stage involves identifying the DG siting decisions in each bus $i$, represented by the binary variable $\delta_{i}^{DG}$, and the corresponding number of DG units $U_{i}^{DG}$ to be installed adhering to the given budget. Constraint (37) ensures the size of installed DG is within the lower limit $\underline{P}^{DG}$ and upper limit $\overline{P}^{DG}$ where $u^{DG}$ is the rating of one DG unit. Constraint (38) ensures that the total cost of DG installation is within the budget $b$ where $c^{DG}_{pu}$ is the per kW cost of DG. Here, the cost of DG is assumed to vary linearly with DG size, and the fixed installation cost is not considered, which can be easily incorporated by just adding the constant term multiplied by the siting variable in the expression. Constraint (39) bounds the number of buses that DG can be installed between $\underline{N}^{DG}\ \text{and}\ \overline{N}^{DG}$. Contrary to prior works in [19] where DG locations were limited to keep the problem size small, we consider all buses as possible candidates for DG, which expands the feasibility space and enhances the quality of the solution.
[…]
$(\textbf{D}\textbf{1}):\quad\min\ \mathbb{E}_{\xi}[Q(x,\xi)]$ (36)
[…]
$\sum_{i=1}^{|\mathcal{B}|}U_{i}^{DG}u^{DG}c^{DG}_{pu}\leq b$ (38)
[…]
$\underline{N}^{DG}\leq\sum_{i=1}^{|\mathcal{B}|}\delta_{i}^{DG}\leq\overline{N}^{DG}$ (39)
III-B Second Stage
The second stage OPF problem minimizes the voltage deviation in each node for a range of scenarios. Any parameter or variable associated with the scenario $\xi\in\Xi$ and phase $\phi\in\Phi$ is superscripted by $\xi$ and $\phi$, respectively. The objective function is given in (40) where $v_{i}$ and $v_{i,ref}$ represent the observed voltage and reference voltage of bus $i$, respectively.
$(\textbf{D}_{\textbf{2}}^{\textbf{v}}):Q(x,\xi)=\min\sum_{i\in\mathcal{B}}\sum_{\phi\in\Phi}|v_{i}^{\phi,\xi}-v_{i,ref}^{\phi}|$ (40)
Since the objective function (40) is non-smooth, it is reformulated as (41) by introducing extra variable $z_{i}^{\phi,\xi}$ and (42), which ensures $z_{i}^{\phi,\xi}$ = $|v_{i}^{\phi,\xi}-v_{i,ref}^{\phi}|$.
$Q(x,\xi)=\min\sum_{i\in\mathcal{B}}\sum_{\phi\in\Phi}z_{i}^{\phi,\xi}$ (41)
$-z_{i}^{\phi,\xi}\leq\ v_{i}^{\phi,\xi}-v_{i,ref}^{\phi}\leq z_{i}^{\phi,\xi}\quad\forall\ i\in\mathcal{B},\quad\forall\ \phi\in\Phi$ (42)
Constraint (43) is the linking constraint between the first and second stages ensuring that the DG power $U_{i}^{DG,\phi,\xi}$ utilized in the given scenario is within the DG capacity decision made in the first stage. Constraints (44)–(47) represent the three-phase unbalanced linearized power flow equations [18]. Active power balance for each bus is maintained by constraint (44) where $P_{i,load}^{\phi}$ denotes base active load of bus $i$ for phase $\phi$. $\epsilon^{\xi}_{i,PV}$ and $\epsilon^{\xi}_{i,load}$ represent uncertainties in PV and load, respectively. $\mathcal{P}_{i}$ and $\mathcal{C}_{i}$ are sets of parent and child buses of $i$ while $P_{ij}^{\phi}$ and $i_{sub}$ denote the active power flowing from bus $i$ to $j$ and substation bus index respectively. Similarly, reactive power balance at the given bus is given by (45) where $Q$ represents the reactive power and $\theta^{DG}$ is the power factor angle of DG. Let $\textbf{v}_{i}^{\xi}$ denote the three-phase voltage square magnitude of bus $i$ while $\textbf{P}_{ij}^{\xi}$ and $\textbf{Q}_{ij}^{\xi}$ are three-phase active and reactive power flowing through line $ij$, respectively. Constraint (46) is the voltage balance equation of line $ij$. $\textbf{r}_{ij}$ and $\textbf{x}_{ij}$ represent the resistance and reactance matrix of line $ij$, respectively. Constraint (47) bounds bus voltage square $\textbf{v}_{i}^{\xi}$ between minimum $\textbf{v}_{\min}$ and maximum $\textbf{v}_{\max}$ allowed values.
$\sum_{\phi\in\Phi}U_{i}^{DG,\phi,\xi}\leq U_{i}^{DG}u^{DG}\quad\forall\ i\in\mathcal{B}$ (43)
[…]
$\begin{gathered}U_{i}^{DG,\phi,\xi}\epsilon_{i,PV}^{\xi}-P_{i,load}^{\phi}\epsilon_{i,load}^{\xi}=\sum_{j\in\mathcal{C}_{i}}P_{ij}^{\phi,\xi}-\sum_{j\in\mathcal{P}_{i}}P_{ji}^{\phi,\xi}\\ \quad\forall\ i\in\mathcal{B}\backslash\{i_{sub}\},\quad\forall\ \phi\in\Phi\end{gathered}$ (44)
$\begin{gathered}U_{i}^{\phi,\xi}\epsilon_{i,PV}^{\xi}\tan\theta^{DG}-Q_{i,load}^{\phi,\xi}\epsilon_{i,load}^{\xi}=\sum_{j\in\mathcal{C}_{i}}Q_{ij}^{\phi,\xi}-\\ \sum_{j\in\mathcal{P}_{i}}Q_{ji}^{\phi,\xi}\quad\forall\ i\in\mathcal{B}\backslash\{i_{sub}\},\quad\forall\ \phi\in\Phi\end{gathered}$ (45)
[…]
$\textbf{v}_{i}^{\xi}-\textbf{v}_{j}^{\xi}=2(\tilde{\textbf{r}}_{ij}\textbf{P}_{ij}^{\xi}+\tilde{\textbf{x}}_{ij}\textbf{Q}_{ij}^{\xi})\quad\forall\ ij\in\mathcal{E}$ (46)
[…]
$\tilde{\textbf{r}}_{ij}={\rm Real}\{\alpha\alpha^{H}\}\otimes\textbf{r}_{ij}+{\rm Im}\{\alpha\alpha^{H}\}\otimes\textbf{x}_{ij}$
[…]
$\tilde{\textbf{x}}_{ij}={\rm Real}\{\alpha\alpha^{H}\}\otimes\textbf{x}_{ij}+{\rm Im}\{\alpha\alpha^{H}\}\otimes\textbf{r}_{ij}$
[…]
$\alpha=[1\;e^{-j2\pi/3}\;e^{j2\pi/3}]^{T}$
[…]
$\textbf{v}_{\min}\leq\textbf{v}_{i}^{\xi}\leq\textbf{v}_{\max}\quad\forall\ i\in\mathcal{B}$ (47)
Constraint (48) ensures that power flowing through line $ij$ is within its thermal rating $\textbf{S}^{rated}_{ij}$. This quadratic constraint is substituted with linear constraints (49), (50) and (51) using polygon based approximations [19].
[…]
$\textbf{P}_{ij}^{\xi^{2}}+\textbf{Q}_{ij}^{\xi^{2}}\leq(\textbf{S}_{ij}^{rated})^{2}\quad\forall\ ij\in\mathcal{E}$ (48)
[…]
$-\sqrt{3}(\textbf{P}_{ij}^{\xi}+\textbf{S}_{ij})\leq\textbf{Q}_{ij}^{\xi}\leq-\sqrt{3}(\textbf{P}_{ij}^{\xi}-\textbf{S}_{ij})\quad\forall\ ij\in\mathcal{E}$ (49)
[…]
$-\frac{\sqrt{3}}{2}\textbf{S}_{ij}\leq\textbf{Q}_{ij}^{\xi}\leq\frac{\sqrt{3}}{2}\textbf{S}_{ij}\quad\forall\ ij\in\mathcal{E}$ (50)
[…]
$\sqrt{3}(\textbf{P}_{ij}^{\xi}-\textbf{S}_{ij})\leq\textbf{Q}_{ij}^{\xi}\leq\sqrt{3}(\textbf{P}_{ij}^{\xi}+\textbf{S}_{ij})\quad\forall\ ij\in\mathcal{E}$ (51)
[…]
where,
[…]
$\textbf{S}_{ij}=\textbf{S}_{ij}^{rated}\sqrt{\frac{2\pi/6}{\sin(2\pi/6)}}$
Remark 4.
$\textbf{D}_{\textbf{2}}^{\textbf{v}}$ can be adapted to minimize power loss by replacing objective function in (41) with a quadratic approximated power loss objective and removing (42); the resulting $\textbf{D}_{\textbf{2}}^{\textbf{p}}$ is given in (52). Voltage $v_{i}^{\phi,\xi}$ in the objective term is assumed to be nominal voltage $v_{i,nom}^{\phi}$ [20]. We also analyzed SPAR-OPF with $\textbf{D}_{\textbf{2}}^{\textbf{p}}$, which incorporates a quadratic function in the second stage.
[…]
$(\textbf{D}_{\textbf{2}}^{\textbf{p}}):\quad Q(x,\xi)=\min\sum_{ij\in\mathcal{E}}\sum_{\phi\in\Phi}\frac{(P_{ij}^{\phi,\xi})^{2}+(Q_{ij}^{\phi,\xi})^{2}}{v_{i}^{\phi,\xi}}r_{ij}$ (52)
[…]
III-C Load and PV Uncertainty Model
We consider the temporal relationship between load and PV outputs based on historical observations of load and PV profiles at a fixed time resolution. For the PV output model, we leverage the PVwatts tool developed by the National Renewable Energy Laboratory (NREL) [21]. The irradiance and temperature of the location are considered to obtain a normalized PV output, also called PV multiplier. Since our main focus is not on scenario generation and reduction, a simple but realistic model motivated by stratified and importance sampling is used in this work [22].
Algorithm 2 Scenario Generation and Reduction Algorithm
1:  Input: $\hat{\Xi},N_{t}$
2:  Output: $\Xi=\{\xi_{1},\xi_{2},...,\xi_{N_{t}}\}$
3:  $N_{D}^{\rm agg}\leftarrow\frac{|\hat{\Xi}|}{N_{t}}$
4:  initialize $s=1,b=1,e=N_{D}^{\rm agg}$
5:  while $e\leq|\hat{\Xi}|$ do
6:     for $h=1:24$ do
7:        $\alpha^{\rm load}(s)=\mathbb{E}[\alpha_{h}^{\rm load}(d)]_{d=b}^{e}$$\alpha^{PV}(s)=\mathbb{E}[\alpha_{h}^{PV}(d)]_{d=b}^{e}$
8:        $s\leftarrow s+1$
9:     end for
10:     $b\leftarrow b+N_{D}^{\rm agg},\quad e\leftarrow e+N_{D}^{\rm agg}$
11:  end while
12:  for $s=1:N_{t}$ do
13:     $\alpha_{i}^{\rm load}(s)=\alpha^{\rm load}(s)+\delta({\alpha_{i}^{\rm load}}(s))\quad\forall\ i\in\mathcal{B}$
14:     $\alpha_{i}^{PV}(s)=\alpha^{\rm load}(s)+\delta({\alpha_{i}^{PV}}(s))\ \quad\forall\ i\in\mathcal{B}$
15:  end for
16:  Return: $\Xi\ \text{where}\ \xi_{s}=[\alpha_{i}^{\rm load}(s),\alpha_{i}^{PV}(s)\quad\forall\ i\in\mathcal{B}]$
The algorithm for scenario generation and reduction is presented in Algorithm 2, which takes $\hat{\Xi}$ scenarios consisting of hourly resolution historical normalized load and PV output and provides $N_{t}$ output scenarios. The scenario index of scenario $\xi_{s}\in\Xi$ is denoted by $s$ in the algorithm. The number of days to be aggregated $N_{D}^{\rm agg}$ in a stratum is calculated based on the total historical data available and the number of scenarios desired $N_{t}$. Then, aggregation is done for each hour for these aggregated days, the index of which starts from $b$ to $e$, to obtain scenarios. $\alpha_{h}^{\rm load}(d)\in\hat{\Xi}$ and $\alpha_{h}^{PV}(d)\in\hat{\Xi}$ denote the normalized load and PV multiplier at hour $h$ of day $d$, respectively. $\mathbb{E}[\alpha_{h}^{\rm load}(d)]_{d=b}^{e}$ and $\mathbb{E}[\alpha_{h}^{PV}(d)]_{d=b}^{e}$ denote the expectation of load and PV multiplier, respectively, corresponding to hour $h$ considering days indexed from $b$ to $e$. Then, once all 24 hours are done, the day’s index will be sided by $N_{D}^{\rm agg}$ to obtain scenarios considering the next strata of aggregated days and so on until the last data point in $\hat{\Xi}$ is reached.
This gives the uniform load $\alpha^{\rm load}(s)$, and PV $\alpha^{PV}(s)$, multiplier profile for each scenario $\xi_{s}$ for the given system. In reality, certain variations in load and PV profile exist among buses. Thus, the load and PV multiplier for each bus is made different as shown in steps 13 and 14, respectively, where $\delta({\alpha_{i}^{\rm load}}(s))$ and $\delta({\alpha_{i}^{PV}}(s))$ represent the small deviations from the assumptions of uniform shape for load and PV, respectively, for bus $i$ for scenario $\xi_{s}$. These deviations are generated from Gaussian random noise. Each process is assumed to be independent of the other and identically distributed. The user can choose the standard deviation value depending on the diversity of historical profiles. Finally, the algorithm provides a scenario set $\Xi$, where each scenario contains load and PV multiplier corresponding to each bus.
III-D SPAR-OPF algorithm for DG siting and sizing
Algorithm 3 shows the overall algorithm using SPAR-OPF for DG siting and sizing problem which takes the power flow model, scenarios set, and other required parameters222parameters that are used in modeling, see Section III-A and III-B as input and outputs the planning decisions: DG size $U_{i}^{DG^{\text{final}}}$ and siting decisions $\delta^{DG^{\text{final}}}$. The number of coordinate functions to be learned equals the number of buses, as DG decisions for each bus are linked to the second stage and we consider all buses as potential locations. Consider L as the maximum number of DG units that can be installed in a bus. Let $m_{i}^{l,(k)}\in m_{i}^{(k)}$ denote the slope of the coordinate function $g_{i}$ at point $l\in\mathcal{L}=[0,L]$ during iteration $k$ where $m_{i}^{(k)}$ is its slope vector. The slope $m_{i}^{l,(1)}$ of all coordinate functions is initialized to zero. The maximum number of iterations is set as $k^{\max}$. Then, at iteration $k$, with $m_{i}^{(k)}$ and using (12), the value of $g_{i}$ is updated. $g_{i}^{l,(k)}$ denotes its value at $l$ in the $k^{th}$ iteration. The first stage problem is then solved using either $\textbf{L}_{1}$ or $\textbf{E}_{1}$ and DG decisions (linking variables) $U_{i}^{DG^{(k)}}$ are passed to the second stage. Scenario $\xi$ is selected randomly from the scenario set $\Xi$ and uncertain parameters specified in (44) and (45) in the second stage problem III-B are updated. Then, the second stage problem is solved, and the dual information $\gamma_{i}^{(k)}$ of the linking constraints (43) is obtained. The slope of each coordinate function at the given first stage decision $U_{i}^{DG^{(k)}}$ is then adjusted using $\gamma_{i}^{(k)}$ where updated step size $\alpha^{(k)}$ is utilized. Using the projection algorithm as described in Section II-D, updated slope vector $n_{i}^{(k)}$ is then projected to the convex set to obtain $m_{i}^{(k+1)}$, which is utilized in the next iteration. The process continuous until the convergence criteria are met or the maximum iteration limit is reached.
Algorithm 3 SPAR-OPF for DG Siting and Sizing
1:  Input: $G,\Xi,k^{\max},\mathcal{L},\underline{P}^{DG},\overline{P}^{DG},c^{DG}_{pu},u^{DG},b,\theta^{DG},\textbf{x}_{ij},\textbf{r}_{ij}$
2:  Output: $U_{i}^{DG^{\text{final}}},\delta_{i}^{DG^{\text{final}}}$
3:  $m_{i}^{l,(1)}\leftarrow 0\quad\forall\ i\in\mathcal{B},\ \forall\ l\in\mathcal{L}$
4:  for $k=1,2,\dots,k^{\max}$ do
5:     Update $g_{i}^{l,(k)}\ \text{using}\ m_{i}^{(k)}$
6:     $U_{i}^{DG^{(k)}}\leftarrow$ Solve first stage problem (III-A)
7:     Draw random scenario $\xi\in\Xi$
8:     for each bus $i\in\mathcal{B}$ do
9:        update $\epsilon_{i,PV}^{\xi},\epsilon_{i,load}^{\xi}$ in (44) and (45)
10:        update $U_{i}^{DG}\leftarrow U_{i}^{DG^{(k)}}$ in (43)
11:     end for
12:     Solve second stage problem (III-B)
[…]
14:     Update $\alpha^{(k)}$
15:     for each $\text{each bus}\ i\in\mathcal{B}$ do
16:        for $l=0,1,2,\dots,L$ do
17:           if $l=U_{i}^{DG^{(k)}}$ then
18:              $n_{i}^{l,(k)}=(1-\alpha^{(k)})m_{i}^{l,(k)}+\alpha^{(k)}\gamma_{i}^{(k)}$
19:           else
20:              $n_{i}^{l,(k)}=m_{i}^{l,(k)}$
21:           end if
[…]
23:        end for
24:     end for
25:  end for
26:  Check the termination criteria
27:  Return: $U_{i}^{DG^{\text{final}}}\leftarrow U_{i}^{DG^{(k)}},\delta_{i}^{DG^{\text{final}}}\leftarrow\delta_{i}^{DG^{(k)}}$
IV Results and Discussions
The effectiveness and scalability of SPAR-OPF are illustrated using two test systems: IEEE 123-bus and a 9500-node (modified IEEE 8500-node) distribution test system. We compare solutions from SPAR-OPF with solutions from EF, which provides a measure of optimality. The solution obtained from the proposed framework is also compared with that from PH, which uses scenario-based decomposition. The heuristic parameter $\rho$ for PH is set to 1. We refer the reader to [11] for additional discussion on PH.
The default learning step size rule for SPAR-OPF, which we denote here as “step rule 1” is taken as $\alpha^{(k)}=20/(20+k)$ as suggested by [15]. Furthermore, different step size rules are tested to measure the robustness of the framework with respect to step size selection. Specifically, we define “step size rule 2” as $\alpha^{(k)}=1/k$, and “step size rule 3” as $\alpha^{(k)}=min(1,20/k)$ to cover different varieties. $k^{\max}$ and $\epsilon-radius$ are set to 100 and $10^{-4}$ for termination. The results, including the effects of increasing number of scenarios and system size, are then presented along with statistical validation. Finally, the computational comparison between the two formulations $\textbf{L}_{1}$ and $\textbf{E}_{1}$ is made. In the remainder of this paper, $\textbf{D}_{2}^{\textbf{v}}$ refers to two-stage stochastic optimization problem with voltage deviation minimization as an objective in the second stage, while $\textbf{D}_{2}^{\textbf{p}}$ refers to the one with power loss minimization as an objective.
IV-A Simulation Set-Up
We use Pyomo [13] to model and Gurobi [23] to solve the optimization problems [^1] All the codes implemented to develop SPAR-OPF will be made publicly available once the paper is accepted.. The EF and PH algorithms are implemented using PySP package in Pyomo. All simulations are carried out on a workstation with 32 GB RAM and an Intel Core i9-10900X CPU @ 3.70GHz.
For simulations, we consider $\overline{N}^{DG}=10$, $\underline{P}^{DG}=33\text{kW}$, $\overline{P}^{DG}=333\text{kW}$, $u^{DG}=2\text{kW}$, and $c_{pu}^{DG}=\$1.01$ [24]. A total budget $b$ of \$1500k is assumed. Note that all these parameters can be easily relaxed based on the user’s preference and available resources. All results are generated by averaging the 25 independent random seeds. Unless otherwise mentioned, SPAR-OPF utilized the epigraph $\textbf{E}_{1}$ formulation.
Scenarios are generated and reduced based on Algorithm 2 described in section III-C. We add gaussian noise with zero mean and 10% standard deviation on load and PV profile to obtain bus profiles. Hourly resolution weather data is obtained from the National Solar Radiation Data Base (NSRDB) considering El Paso, the western area of Texas, which is served by the Electric Reliability Council of Texas (ERCOT) for the year 2023 [25]. Hourly resolution historical load data for the year 2023 of the same area is taken from ERCOT [26]. This gives 8760 scenarios representing each hour of the year. Then, 96 and 1200 scenarios are generated for each system using Algorithm 2. For instance, Fig. 2(a) and Fig. 2(b) show the load and PV profile of 96 scenarios for IEEE 123-bus system, respectively, where solid line denotes the common shape and shaded curves are profiles for each bus.
(a) Bus load profile
(b) Bus PV profile
Figure 2: Bus load and PV profile for IEEE 123 bus system.
IV-B Illustration of SPAR-OPF
In this section, we illustrate the performance of SPAR-OPF on the IEEE 123 bus system with 96 and 1200 scenarios. The total load of this system is 3431.3 kW and 1931.9 kvar. A system base of 1 MVA and 4.16 kV is considered.
IV-B1 Case I: With 96 scenarios
First, we present a visualization that illustrates the learning of the coordinate functions by SPAR-OPF. Recall that each coordinate function represents value as a function of the corresponding linking variable. Here, we learn a total of 123 coordinate functions, as DG sizing decisions (corresponding to each bus) are linked to the second stage. Fig. 3 shows 9 such coordinate functions for the first five iterations with dark color showing the later iterations for problem $\textbf{D}_{2}^{\textbf{v}}$. The x-axis shows the decision points, and the y-axis shows the value of the function. It can be observed that in every iteration, the function updates its value due to changes in slope information from the algorithm. For the coordinate function $g_{1}$, the function remains constant at every decision point, i.e., equal to the initialized zero value. This function is associated with the substation bus, and installing DG at the substation bus does not add any value to the objective function considered here.
Figure 3: Illustration of coordinate functions learning. The $x$-axis shows the decision points and the $y$-axis shows the value of the coordinate function.
Fig. 4(a) shows the quality of function approximations with respect to iterations for $\textbf{D}_{2}^{\textbf{v}}$. It can be seen that all step size rules give satisfactory performance with a quality of more than 98%. The solution quality is consistent for $\textbf{D}_{2}^{\textbf{p}}$ as well, as shown in Fig. 4(b).
(a) $\textbf{D}_{\textbf{2}}^{\textbf{v}}$
(b) $\textbf{D}_{\textbf{2}}^{\textbf{p}}$
Figure 4: Quality of function approximation with SPAR-OPF.
Table I presents the results from the SPAR-OPF, EF, and PH where columns “Obj” and “Time” indicate the objective value and time taken to solve the problem, respectively. It is important to mention that all algorithms were implemented on same computing environment. For $\textbf{D}_{2}^{\textbf{v}}$, it is observed that the SPAR-OPF objective value is very close to the true objective value (EF column) with a gap of 0.09 (0.44%). The last row shows the solutions for $\textbf{D}_{2}^{\textbf{p}}$ where the solution obtained from the SPAR-OPF is very close to that of the optimal value from EF with a gap of 0.6 kW (0.93%). Note that the proposed framework gives a better solution than PH for $\textbf{D}_{2}^{\textbf{v}}$, whereas later performs better for $\textbf{D}_{2}^{\textbf{p}}$ for the considered problem; nevertheless, both solutions are close to the respective optimal values.
TABLE I: IEEE 123 system with 96 scenarios
Problem SPAR-OPF EF PH
Obj Time(s) Obj Time(s) Obj Time(s)
$\textbf{D}_{\textbf{2}}^{\textbf{v}}$ 20.32 53.6 20.23 47.94 21.98 537
$\textbf{D}_{\textbf{2}}^{\textbf{p}}$ 64.7 kW 58.8 64.1 kW 686 64.24 kW 6767
Note: The objective values for $\textbf{D}_{\textbf{2}}^{\textbf{v}}$ are unitless.
SPAR-OPF and EF took comparable time to solve $\textbf{D}_{2}^{\textbf{v}}$, but PH took longer to solve. For the quadratic objective problem $\textbf{D}_{2}^{\textbf{p}}$, EF took 11.6x more time than SPAR-OPF. Due to the quadratic objective in the second stage, EF involves solving a huge MINLP problem, whereas SPAR-OPF solves smaller MILP in the first stage and QP in the second stage, thereby decreasing the overall solve time. Due to this reason, SPAR-OPF took comparable time to solve $\textbf{D}_{2}^{\textbf{v}}$ and $\textbf{D}_{2}^{\textbf{p}}$, whereas time for EF increases by 14x to solve the later. This is another advantage of the proposed framework, which decomposes the MINLP problem into easier sub-problems. PH takes a comparatively longer time to solve both problems. This performance time might get lower if PH had been implemented in a high-performance computing device. However, PH also involves MINLP subproblems, which would take more time even if scenario subproblems are solved in parallel.
IV-B2 Case II: with 1200 scenarios
In this section, we illustrate statistical bounds and compare solution quality with EF and PH using a significantly larger number of scenarios, and analyze computational time. Table II presents the results for $\textbf{D}_{\textbf{2}}^{\textbf{v}}$ considering 1200 scenarios. It is observed that the extensive formulation for this case has 5.28 millions rows and 2.33 millions columns whereas case I has 0.42 millions rows and 0.186 millions columns, which illustrates the increasing complexity of the problem with an increase in the number of scenarios. The column $N_{s}$ in Table II indicates the batch size taken for each random seed while solving the problem using SPAR-OPF. The columns “Obj”, “LB CI”, and “UB CI” refer to objective value, lower bounds confidence interval and upper bound confidence interval for SPAR-OPF respectively. Similarly, column “BG” refers to the bounds gap with $(1-2\alpha)$, where $\alpha=0.05$, confidence level calculated based on the difference between the upper value of the upper bound confidence interval and the lower value of the lower bound confidence interval. The number within the bracket shows the percentage bounds gap to the objective value. Note that the objective value from SPAR-OPF lies within the lower bound and upper bound confidence interval. For smaller batch sizes, the width of the confidence interval and bounds gap are higher as the batch size is insufficient to capture the uncertainty set. The proposed framework obtains a solution closer, and better than the PH solution, to the optimal solution obtained from the EF. The width of UB CI does not change that much as all scenarios are simulated, and less variability is observed.
It is important to mention that a bounds gap of $x\%$ does not necessarily mean that the objective value of SPAR-OPF is $x\%$ away from the optimal solution. It measures the tightness of bounds at the current solution. Take an instance, for case with $N_{s}=200$, the objective value from SPAR-OPF is only 0.05%, far from the optimal value though the BGis 2.1%. It is observed that the proposed framework took 67.8 seconds per seed, whereas EF and PH took 1.86 hours and 2.89 hours to solve the problem, respectively.
Table III presents results for problem $\textbf{D}_{\textbf{2}}^{\textbf{p}}$ with varying batch size where bounds gap and optimality gap are less than 4% and 0.62% respectively for batch size $N_{s}$ beyond 150. EF, which involved solving a large MINLP problem, took 16.8 hours to complete, while PH, requiring the solution of 1200 MINLP problems per iteration, took 22.4 hours. On the other hand, SPAR-OPF took 79.6 seconds per seed to solve the same problem. This shows the benefit of SPAR-OPF, which takes one scenario at a time, making it practical even for environments with limited resources.
TABLE II: IEEE 123 system with 1200 scenarios for $\textbf{D}_{\textbf{2}}^{\textbf{v}}$
$N_{s}$ Obj LB CI UB CI BG (%) EF PH
50 16.99 [16.67-17.31] [17.41-17.53] 0.86 (5.1) 17.31 19.25
100 17.09 [16.87-17.32] [17.3-17.45] 0.58 (3.4)
150 17.25 [17.08-17.42] [17.41-17.52] 0.44 (2.5)
200 17.3 [17.13-17.48] [17.4-17.5] 0.37 (2.1)
Note: All Obj, LB CI, UB CI and BG values are unitless.
TABLE III: IEEE 123 system with 1200 scenarios for $\textbf{D}_{\textbf{2}}^{\textbf{p}}$
$N_{s}$ Obj LB CI UB CI BG (%) EF PH
50 48.53 [47.04-50.02] [50.26-50.62] 3.58 (7.4) 49.39 49.46
100 49.02 [47.95-50.07] [50.21-50.41] 2.46 (5)
150 49.7 [48.9-50.52] [50.33-50.83] 1.93 (3.8)
200 49.62 [48.78-50.48] [50.21-50.54] 1.76 (3.5)
Note: All Obj, LB CI, UB CI and BG values are in kW.
IV-C Scalability assessment with system size
In this section, we test the framework’s scalability considering the 9500-node system, a modified IEEE 8500-node test system, with a total load of 13668 kW + 3765 kvar [27]. All secondary loads were referred to the primary side of the transformer. A system base of 1 MVA and 12.47 kV is assumed. We consider 1200 scenarios. EF would result in 112 million rows and 49.4 million columns, which is significantly challenging for any commercial solver to solve this optimization problem as a single large problem in polynomial time. Thus, EF and PH could not solve this problem as it became intractable using the considered workstation. Table IV and Table V present the results for the problem $\textbf{D}_{\textbf{2}}^{\textbf{v}}$ and $\textbf{D}_{\textbf{2}}^{\textbf{p}}$ respectively from SPAR-OPF. The obtained solutions are well within the tight bounds where the bounds gap is less than 3% for the batch size $N_{s}\geq 150$. SPAR-OPF took 5.25 minutes and 5.5 minutes per seed to solve $\textbf{D}_{\textbf{2}}^{\textbf{v}}$ and $\textbf{D}_{\textbf{2}}^{\textbf{p}}$ respectively. This illustrates the scalability of proposed framework, making it suitable even for real-time applications under limited computing resources.
TABLE IV: 9500-node system with 1200 scenarios for $\textbf{D}_{\textbf{2}}^{\textbf{v}}$
$N_{s}$ Obj LB CI UB CI BG (%)
50 622.7 [609.9-635.5] [622.5-639.1] 29.2(4.68)
100 623.9 [612.3-635.6] [624.1-639.9] 27.6 (4.42)
150 627.8 [619.2-636.5] [623.2-637] 17.8 (2.8)
200 627.9 [621.2-634.2] [624.1-638] 16.8 (2.7)
Note: All Obj, LB CI, UB CI and BG values are unitless.
TABLE V: 9500-node system with 1200 scenarios for $\textbf{D}_{\textbf{2}}^{\textbf{p}}$
$N_{s}$ Obj LB CI UB CI BG (%)
50 357 [349.2-364.8] [367.2-370.9] 21.7 (6.1)
100 361.2 [355.5-366.9] [368.2-371.9] 16.4 (4.5)
150 364.1 [360.2-367.9] [367.7-371.1] 10.9 (3.0)
200 364.8 [361.2-368.4] [367.9-371.8] 10.6 (2.9)
Note: All Obj, LB CI, UB CI and BG values are in kW.
IV-D Computational comparison between $\textbf{L}_{\textbf{1}}$ and $\textbf{E}_{\textbf{1}}$
In this section, the computational performance between two formulations obtained using $\textbf{L}_{\textbf{1}}$ and $\textbf{E}_{\textbf{1}}$ is made. Table VI shows the computational performance of these two formulations for different test cases and objectives. The values corresponding to columns $\textbf{D}_{\textbf{2}}^{\textbf{v}}$ and $\textbf{D}_{\textbf{2}}^{\textbf{p}}$ represent the time in seconds per iteration while solving the corresponding problems. The number of rows and columns in the optimization problem for $\textbf{D}_{\textbf{2}}^{\textbf{v}}$ is slightly higher due to additional constraint (42) and number for $\textbf{D}_{\textbf{2}}^{\textbf{v}}$ is reported. Leveraging the convexity of the value function and using only a few additional continuous variables, $\textbf{E}_{\textbf{1}}$ is observed to perform better (4x less time) than $\textbf{L}_{\textbf{1}}$. Even for the larger system size, $\textbf{E}_{\textbf{1}}$ took only 5 seconds per iteration to solve the problem. Based on all simulations, 80 iterations were required on average for the convergence. This confirms that the proposed framework applies to realistic large-scale power systems.
TABLE VI: Computational comparison between $\textbf{L}_{\textbf{1}}$ and $\textbf{E}_{\textbf{1}}$
System Method $\textbf{D}_{\textbf{2}}^{\textbf{v}}(s)$ $\textbf{D}_{\textbf{2}}^{\textbf{p}}(s)$ Rows Columns
IEEE 123 bus system $\textbf{L}_{\textbf{1}}$ 2.08 2.1 524 21.9k(21.8k binary)
$\textbf{E}_{\textbf{1}}$ 0.67 0.69 21.9k 391 (130 binary)
9500-node system $\textbf{L}_{\textbf{1}}$ 20.1 20.3 10.9k 463.7k (414.3k binary)
$\textbf{E}_{\textbf{1}}$ 5 5.03 463.7k 8.2k (2.7k binary)
V Conclusion and Future Work
We developed the SPAR-OPF framework to solve OPF-based two-stage stochastic optimization problems in power systems. To the best of our knowledge, this is the first to solve OPF-based two-stage stochastic optimization problems using separable projective approximations of value function. We provided two formulations for the learned functions and a statistical validation procedure. The proposed framework is illustrated on two standard three-phase unbalanced distribution system test cases considering DG siting and sizing problem. Results illustrated that SPAR-OPF provides near-optimal solutions with significantly lower solve times as compared to state-of-the-art methods. The scalability of the proposed framework is validated with a larger system size and a higher number of scenarios, making it suitable even for resource-constrained environments. As a part of future work, we will explore heuristics such as MIP relaxation, exploration vs. exploitation, and successive MIP gap reduction to further reduce the solution time without affecting the final solution. The proposed framework will further be demonstrated on classical stochastic problems in power systems such as unit commitment and dynamic economic dispatch.